# Multi-Agent Guided Policy Search for Non-Cooperative Dynamic Games

### v1

@@ I Introduction
These limitations highlight the complementary strengths of MARL and mode-based optimization, suggesting that a hybrid method could provide both scalability and stability. In this paper, we move beyond prior efforts that integrate such approaches for cooperative multi-agent control [10, 11] and ask: how can we integrate these two approaches to improve policy learning for non-cooperative dynamic games?
@@
We propose a framework that incorporates this prior into the reward function to regularize MA-PG updates. In doing so, it offers a new mechanism for stabilizing MARL that outperforms existing techniques based on entropy-driven exploration [12, 13]. Our method not only achieves faster convergence but also yields neural network policies that can be executed in real time for complex multi-agent dynamic games, where classical model-based optimization methods are computationally demanding for timely decision-making. Our contributions are:
@@ II Related works
Model-based priors have been explored to reduce variance in RL [40]. Unlike [40], which regularizes the policy output directly, our method incorporates the regularization into the reward function. Prior works such as [12, 13, 41] stabilize MA-PG in finite-horizon games using entropy-driven exploration. In contrast, we show that suitable policy guidance ensures stability in infinite-horizon settings, which could be more practical for real-time decision-making since the optimal policy is time-invariant and avoids computing complex time-varying Nash equilibrium policies.
@@ III-A Problem settings
${\pi}\leftarrow{\pi}-\eta\nabla w({\pi}).$ (7)
III-B Non-cooperative LQ games
In this subsection, we introduce a special class of non-cooperative games, LQ games [1], in order to present a relatively simple case where we can theoretically characterize the performance of policy gradient methods. In infinite-horizon LQ games, we consider
$\min_{{\pi}^{i}}\$$\mathbb{E}_{x_{0}\sim{\mathcal{D}}}\Big[\sum_{{t}=0}^{\infty}{x}_{t}^{\top}{Q}^{i}{x}_{t}+{u}_{t}^{\top}{R}^{i}{u}_{t}^{i}\Big]$ (8)
s.t. ${x}_{{t}+1}={A}{x}_{t}+{B}{u}_{t},\;{u}_{t}\sim{\pi}^{i}({x}_{t}),\;\forall t=0,1,\dots\vskip-2.5pt$
where discount factor is set to 1. Note that for policies resulting in stable closed-loop dynamics, we do not need to introduce the discount. More formally, we consider the following assumption:
@@ IV Analysis: Stability and bias of the guided policy optimization in LQ games
    As we increase $\rho$, the regularized Nash equilibrium $\hat{K}^{*}$ will exhibit locally stable policy gradient dynamics, with $\lambda_{\min}>0$.
@@
    where $\bar{A}_{\hat{K}}=A-\sum_{i=1}^{N}B_{i}\hat{K}_{i}$, and $\hat{P}_{\hat{K}}^{i}$ is given as the unique positive definite solution of $\hat{P}_{\hat{K}}^{i}={\bar{A}_{\hat{K}}}^{\top}\hat{P}_{\hat{K}}^{i}\bar{A}_{\hat{K}}+\hat{K}^{i}{}^{\top}R^{i}\hat{K}^{i}+Q^{i}+\rho{(\hat{K}^{i}-\check{K}^{i})}^{\top}R^{i}(\hat{K}^{i}-\check{K}^{i})$.
@@
Figure 1: Policy guidance stabilizes infinite-horizon LQ game dynamics, enables convergence to a close neighborhood of the true Nash equilibrium—even if the guidance is wrong. The first two rows are the trajectories of the values of the two entries of $K^{1},K^{2}$ under policy gradient with different values of $\rho$ and a biased guiding policy $\check{K}$, which may come from an inaccurate dynamics model. A small $\rho$ ($\geq 0.001$) is sufficient to stabilize the policy gradient dynamics nearby the ground truth feedback Nash equilibrium $K^{*}$. However, as $\rho$ increases, the bias in the converged policies introduced by the guidance also grows.
Proposition 1 indicates that adding the guidance term from (13) with a sufficiently large $\rho$ stabilizes the policy gradient dynamics, guaranteeing the local exponential convergence of policy gradient methods to a stationary point even if the original policy gradient dynamics is unstable (i.e., $\lambda_{\min}<0$). Denote by $R\in\mathbb{R}^{m\times m}$ a diagonal block matrix where the $i$-th diagonal block is composed by $R^{i}$. This stability is primarily achieved because the term $(\rho\cdot R)$ shifts all eigenvalues of the pseudo-gradient’s Jacobian $(\nabla w(K)+\rho\cdot R)$ to the positive domain. However, the distance (bias) between the converged solution and the true Nash equilibrium depends jointly on the problem dynamics, $\rho$, and the quality of the guidance policy. We include a numerical study in Figure 1 to support this. Specifically, there are three cases:
@@
    Further increasing $\rho$, we will enhance stability but the converged solution may deviate from the Nash equilibrium of the original game and biases more towards the guidance policy.
@@ V-A Efficient computation of local guidance for MA-PG
To achieve this, we introduce the time-varying finite-horizon linear quadratic games, with a given initial state ${\bar{x}}_{0}$:
@@
s.t. ${x}_{{t}+1}={A}_{t}{x}_{t}+\sum_{i=1}^{N}{B}_{t}^{i}{u}_{t}^{i},\ {x}_{0}={\bar{x}}_{0},\ t\in[T]$
[…]
@@
${Q}_{t}^{i}=\frac{1}{2}\nabla^{2}_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $q_{t}^{i}=\frac{1}{2}\nabla_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$
${R}_{t}^{i}=\frac{1}{2}\nabla^{2}_{{\bar{u}}^{i}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $r_{t}^{i}=\frac{1}{2}\nabla_{{\bar{u}}_{t}^{i}}{c}^{i}({\bar{x}}_{x},{\bar{u}}_{t}).$
@@
$\hat{{c}}^{i}({x},u^{1},\dots,u^{N}):={c}^{i}({x},u^{1},\cdots,u^{N})+\rho\|u^{i}-\check{{\pi}}^{i}(x,{\pi_{\theta}})\|_{2}^{2}\vskip-3.00003pt$ (16)
@@ 2 Set data buffer $D=\{\}$
6      For each $j\in\{1,\dots,M_{r}\}$, simulate trajectories $\tau^{(j)}=\{x_{0}^{(j)}\!,u_{0}^{(j)}\!,\dots,x_{T_{r}+1}^{(j)}\}$, under the policy $\{{\pi_{\theta}}^{i}\}_{i=1}^{N}$.
7      Update data buffer, $D\leftarrow\{\tau^{(j)}\}_{j=1}^{M_{r}}$
@@ 9         Sampling a batch of states $\{x^{(j)}\}_{j=1}^{M_{b}}$
10         Obtain the guidance $\{\check{u}^{(j)}\}_{j=1}^{M_{b}}$ for each $\{x^{(j)}\}_{j=1}^{M_{b}}$, as outlined in Section V-A
[…]
12        Optimizing value function via minimizing Bellman error [34]:
13       $\theta^{\prime}\leftarrow\theta^{\prime}-{\eta}\nabla\sum_{j=1}^{M_{b}}\|V_{\theta^{\prime}}^{i,\pi_{\theta}}(x^{(j)})-c^{i}(x^{(j)},\pi_{\theta}(x^{(j)}))-\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}(f(x^{(j)},\pi_{\theta}(x^{(j)})))\|_{2}$
14
15
@@ VI Experiments
1) LQ game. As noted in [5], MA-PG methods in LQ games exhibit limit-cycle behaviors and fail to converge to Nash equilibria when the Jacobian of $w(K)$ in (9) has both positive and negative eigenvalues. We reproduce this phenomenon and study the stabilizing effect of our guidance regularization term. Using the system dynamics described in Appendix VIII-A, we solve for the ground truth Nash equilibrium policies $K^{1*},K^{2*}$ via the Lyapunov iteration from [5]. We introduce an artificial bias to the guiding policies $\check{K}^{1},\check{K}^{2}$ in the regularized cost (13). The resulting policy trajectories under varying regularization weights $\rho$ are shown in Figure 1.
@@
3) Six-player strategic basketball formation.
We demonstrate that MA-GPS scales to larger numbers of agents and converges faster to high-performing policies than IPPO without guidance, measured in training time. To illustrate, we simulate a simplified basketball formation where both offense and defense teams consist of three agents: a shooting guard, a center, and a forward. Defensive agents aim to stay between their assigned opponent and the basket. On offense, the center positions near the basket ($r=1.5$), the shooting guard holds the three-point line ($r=4$) to prepare for shots, and the forward sets a screen by moving toward the defender covering the shooting guard.
The agents’ objectives include mild non-differentiabilities. For example, letting $p_{y}$ denote a player’s y-position, a term like $\max(p_{y},0)$ imposes a one-sided penalty with a kink at the baseline ($p_{y}=0$), discouraging players from crossing it. This illustrates that MA-GPS can handle objectives beyond perfectly smooth settings, broadening its applicability to a wider range of multi-agent tasks.
@@
We measure wall-clock computation time $\tau$ for iLQGames with horizon length $T=20$ on two tasks: three-vehicle platooning ($\tau=0.0485\pm 0.0042\,\mathrm{s}$) and six-player basketball formation ($\tau=46.81\pm 4.18\,\mathrm{s}$). Because iLQGames uses finite-horizon formulations, its computational time scales quickly with horizon length, player count, and objective complexity. In contrast, MA-GPS shifts computation to offline training and directly solves the infinite-horizon case, producing real-time executable neural network policies ($\tau=0.0016\pm 0.0012\,\mathrm{s}$) for long-horizon strategic decision-making.

### v3

@@ I Introduction
These limitations highlight the complementary strengths of MARL and model-based optimization, suggesting that a hybrid method could provide both scalability and stability. In this paper, we move beyond prior efforts that integrate such approaches for cooperative multi-agent control [10, 11] and ask: how can we integrate these two approaches to improve policy learning for non-cooperative dynamic games?
@@
We propose a framework that incorporates this prior into the reward function to regularize MA-PG updates. This constitutes a new mechanism for stabilizing MARL, outperforming existing MARL stabilization approaches that primarily rely on entropy-driven exploration [12, 13]. Beyond faster convergence, our method yields neural network policies that can be executed in real time for complex multi-agent dynamic games, where classical model-based optimization methods are often computationally demanding for timely decision-making. Our contributions are:
@@ II Related works
Model-based priors have been explored to reduce variance in RL [40]. Unlike [40], which regularizes the policy output directly, our method incorporates regularization into the reward function. Moreover, prior works such as [12, 13, 41] stabilize MA-PG in finite-horizon games using entropy-driven exploration. In contrast, we show that suitable policy guidance ensures stability in infinite-horizon settings, which could be more practical for real-time decision-making since the optimal policy is time-invariant and avoids computing complex time-varying Nash equilibrium policies.
@@ III-A Problem settings
${\pi}\leftarrow{\pi}-\eta w({\pi}).$ (7)
III-B Non-cooperative linear quadratic games
In this subsection, we introduce a special class of non-cooperative games known as linear–quadratic (LQ) games [1], a simplified setting that allows theoretical characterization of the performance of policy gradient methods. In infinite-horizon LQ games, we consider for each agent $i\in[N]$,
$\min_{{\pi}^{i}}\$$\mathbb{E}_{x_{0}\sim{\mathcal{D}}}\Big[\sum_{{t}=0}^{\infty}{x}_{t}^{\top}{Q}^{i}{x}_{t}+{u}_{t}^{i\top}{R}^{i}{u}_{t}^{i}\Big]$ (8)
s.t. ${x}_{{t}+1}={A}{x}_{t}+\sum_{i=1}^{N}{B}^{i}{u}_{t}^{i},\;{u}_{t}^{i}\sim{\pi}^{i}({x}_{t}),\;\forall t=0,1,\dots$
where $A\in\mathbb{R}^{n\times n}$, $B^{i}\in\mathbb{R}^{n\times m_{i}}$, $Q^{i}\in\mathbb{R}^{n\times n}$ and $R^{i}\in\mathbb{R}^{m_{i}\times m_{i}}$, $\forall i\in[N]$. We consider the following assumption:
@@ IV Analysis: Stability and bias of the guided policy optimization in LQ games
    As $\rho$ increases, the regularized Nash equilibrium $\hat{K}^{*}$ will exhibit locally stable policy gradient dynamics, with all eigenvalues of the regularized pseudo-gradient $(\nabla w(K)+\rho\cdot R)$ having positive real parts.
@@
    where $\bar{A}_{\hat{K}}=A-\sum_{i=1}^{N}B^{i}\hat{K}^{i}$, and $\hat{P}_{\hat{K}}^{i}$ is given as the unique positive definite solution of $\hat{P}_{\hat{K}}^{i}={\bar{A}_{\hat{K}}}^{\top}\hat{P}_{\hat{K}}^{i}\bar{A}_{\hat{K}}+\hat{K}^{i}{}^{\top}R^{i}\hat{K}^{i}+Q^{i}+\rho{(\hat{K}^{i}-\check{K}^{i})}^{\top}R^{i}(\hat{K}^{i}-\check{K}^{i})$.
@@
Figure 1: Policy guidance stabilizes infinite-horizon LQ game dynamics, enables convergence to a close neighborhood of the true Nash equilibrium—even if the guidance is wrong. Each row corresponds to one of $K^{1}$ and $K^{2}$. In each plot, the two axes represent the entries $K^{i,1}$ and $K^{i,2}$ of the corresponding $K^{i}$, and the curves depict their trajectories under policy gradient with different values of $\rho$ and a biased guiding policy $\check{K}$ (potentially derived from an inaccurate dynamics model). A small $\rho$ ($\geq 0.001$) is sufficient to stabilize the policy gradient dynamics near the ground truth feedback Nash equilibrium $K^{*}$. However, as $\rho$ increases, the bias in the converged policies introduced by the guidance also grows.
Proposition 1 indicates that adding the guidance term from (13) with a sufficiently large $\rho$ stabilizes the policy gradient dynamics, guaranteeing the local exponential convergence of policy gradient methods to a stationary point even if the original policy gradient dynamics is unstable (i.e., $\lambda_{\min}<0$). Denote by $R\in\mathbb{R}^{m\times m}$ a diagonal block matrix where the $i$-th diagonal block is composed by $R^{i}$. This stability is achieved because the term $(\rho\cdot R)$ shifts all eigenvalues of the regularized pseudo-gradient’s Jacobian $(\nabla w(K)+\rho\cdot R)$ to the positive domain. However, the distance (bias) between the converged solution and the true Nash equilibrium depends jointly on the problem dynamics, $\rho$, and the quality of the guidance policy. We include a numerical study in Figure 1 to support this. Specifically, there are three cases:
@@
    Further increasing $\rho$, we will enhance stability, but the converged solution may deviate from the Nash equilibrium of the original game and be biased more towards the guidance policy.
@@ V-A Efficient computation of local guidance for MA-PG
To achieve this, we introduce the time-varying finite-horizon linear quadratic games, with a given initial state ${\bar{x}}_{0}$ and $i\in[N]$:
@@
s.t. ${x}_{{t}+1}={A}_{t}{x}_{t}+\sum_{i=1}^{N}{B}_{t}^{i}{u}_{t}^{i},\ {x}_{0}={\bar{x}}_{0},$
$\forall t=0,\dots,T$
The horizon length $T$ is user-defined and can be relatively short if computational resources are limited. In our experiments (Section VI), we find $T=10$ or $20$ sufficient for generating effective guidance, though the optimal choice depends on game-specific dynamics and costs. We propose training a neural network policy $\pi_{\theta}$} for nonlinear games via model-based guided policy gradients. Below, we describe the procedure for computing the reference trajectory used to obtain policy guidance. Consider a $T$-horizon nominal state trajectory $\{{\bar{x}}_{t}\}_{{t}=0}^{T+1}$ steered by the neural network control $\{{\bar{u}}_{t}\}_{{t}=0}^{T}$, with ${\bar{u}}_{t}\sim{\pi_{\theta}}({\bar{x}}_{t})$, ${\bar{x}}_{{t}+1}={f}({\bar{x}}_{t},{\bar{u}}_{t}^{1},\dots,{\bar{u}}_{t}^{N}),\ \ \forall t=0,\dots,T$. We linearize the dynamics and quadraticize the cost to derive a local LQ game approximation around the nominal trajectory $\{{\bar{x}}_{0},{\bar{u}}_{0},\cdots,{\bar{x}}_{T+1}\}$, for all $i\in[N]$,
@@
${Q}_{t}^{i}=\nabla^{2}_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $q_{t}^{i}=\nabla_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$
${R}_{t}^{i}=\nabla^{2}_{{\bar{u}}^{i}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $r_{t}^{i}=\nabla_{{\bar{u}}_{t}^{i}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}).$
@@
$\hat{{c}}_{\rho}^{i}({x},u^{1},\dots,u^{N}):={c}^{i}({x},u^{1},\cdots,u^{N})+\rho\|u^{i}-\check{{\pi}}^{i}(x,{\pi_{\theta}})\|_{2}^{2}\vskip-3.00003pt$ (16)
@@ 2 Set data buffer $D=\{\}$
6      For each $j\in\{1,\dots,M_{r}\}$, simulate a trajectory $\xi^{(j)}=\{x_{0}^{(j)}\!,u_{0}^{(j)}\!,\dots,x_{T_{r}+1}^{(j)}\}$, under the policies $\{\pi_{\theta}^{i}\}_{i=1}^{N}$.
7      Update data buffer, $D\leftarrow\{\xi^{(j)}\}_{j=1}^{M_{r}}$
@@ 9         Sampling a batch of states $\{x^{(j)}\}_{j=1}^{M_{b}}$
10         Optimizing policy: $\theta\leftarrow\theta-{\eta}\nabla_{\theta}\sum_{j=1}^{M_{b}}\Big[\hat{{c}}_{\rho_{k}}^{i}\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)+\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}\Big(f\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)\Big)\Big]$
11        Optimizing value function via minimizing Bellman error [34]: $\theta^{\prime}\leftarrow\theta^{\prime}-{\eta}\nabla_{\theta^{\prime}}\sum_{j=1}^{M_{b}}\Big\|V_{\theta^{\prime}}^{i,\pi_{\theta}}(x^{(j)})-\hat{{c}}_{\rho_{k}}^{i}\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)-\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}\Big(f\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)\Big)\Big\|_{2}$
12
13
@@ VI Experiments
1) LQ game. As noted in [5], MA-PG methods in LQ games exhibit limit-cycle behaviors and fail to converge to Nash equilibria when the Jacobian of $w(K)$ in (9) has both positive and negative eigenvalues. We reproduce this phenomenon in an LQ game example, detailed in the Appendix, and analyze the stabilizing effect of the proposed guidance regularization term. We solve for the ground truth Nash equilibrium policies $K^{1*},K^{2*}$ via the Lyapunov iteration from [5]. We introduce an artificial bias to the guiding policies $\check{K}^{1},\check{K}^{2}$ in the regularized cost (13). The resulting policy trajectories under varying regularization weights $\rho$ are shown in Figure 1.
@@
3) Six-player strategic basketball formation. We demonstrate that MA-GPS scales to larger numbers of agents and converges faster to high-performing policies than IPPO without guidance, measured in training time. To illustrate, we simulate a simplified basketball formation where both offense and defense teams consist of three agents: a shooting guard, a center, and a forward. Defensive agents aim to stay between their assigned opponent and the basket. On offense, the center positions near the basket ($r=1.5$), the shooting guard holds the three-point line ($r=4$) to prepare for shots, and the forward sets a screen by moving toward the defender covering the shooting guard.
We note that some agents’ objectives include mild non-differentiabilities. For instance, a player’s objective may contain a term like $\max(p_{y},0)$, where $p_{y}$ denotes that player’s y-position. This one-sided penalty introduces a kink at the baseline ($p_{y}=0$), discouraging the player from crossing it. Our experiments indicate that MA-GPS can accommodate such objectives beyond perfectly smooth settings, thereby broadening its applicability to a wider range of multi-agent tasks.
@@
We measure wall-clock computation time $\tau$ for iLQGames with horizon length $T$ on two tasks: three-vehicle platooning ($T=20,\tau=0.0485\pm 0.0042\,\mathrm{s}$) and six-player basketball formation ($T=100,\tau=0.64\pm 0.15\,\mathrm{s}$). Because iLQGames uses finite-horizon formulations, its computational time scales quickly with horizon length, player count, and objective complexity. In contrast, MA-GPS shifts computation to offline training and directly solves the infinite-horizon case, producing real-time executable neural network policies ($\tau=0.0016\pm 0.0012\,\mathrm{s}$) for long-horizon strategic decision-making.